OpenStudy (anonymous):

Describe, in detail, how to make 3.0 liters of a 0.98 molar HCl solution from a 10.5 molar HCl stock solution.

OpenStudy (australopithecus):

you know, You want, 3.0L 0.98M you have, 10.5M You can use ratios to solve this, \[\frac{Concentration Wanted}{Volume Wanted} = \frac{Stock Concentration}{Aliquot Needed}\] Alternatively, you can simplify this formula by simply multiplying your final volume by a ratio of concentrations to find the volume of your aliquot \[(\frac{StockConcentration}{ConcentrationWanted})VolumeWanted = Aliquot Needed\]